# UI-R1: Enhancing Action Prediction of GUI Agents by Reinforcement Learning

### v1

@@ Abstract
The recent DeepSeek-R1 has showcased the emergence of reasoning capabilities in LLMs through reinforcement learning (RL) with rule-based rewards. Building on this idea, we are the first to explore how rule-based RL can enhance the reasoning capabilities of multimodal large language models (MLLMs) for graphic user interface (GUI) action prediction tasks. To this end, we curate a small yet high-quality dataset of 136 challenging tasks, encompassing five common action types on mobile devices. We also introduce a unified rule-based action reward, enabling model optimization via policy-based algorithms such as Group Relative Policy Optimization (GRPO). Experimental results demonstrate that our proposed data-efficient model, UI-R1-3B, achieves substantial improvements on both in-domain (ID) and out-of-domain (OOD) tasks. Specifically, on the ID benchmark AndroidControl, the action type accuracy improves by 15%, while grounding accuracy increases by 10.3%, compared with the base model (i.e. Qwen2.5-VL-3B). On the OOD GUI grounding benchmark ScreenSpot-Pro, our model surpasses the base model by 6.0% and achieves competitive performance with larger models (e.g., OS-Atlas-7B), which are trained via supervised fine-tuning (SFT) on 76K data. These results underscore the potential of rule-based reinforcement learning to advance GUI understanding and control, paving the way for future research in this domain.
@@ 1 Introduction
Rule-based reinforcement learning (RL) or reinforcement fine-tuning (RFT) has recently emerged as an efficient and scalable alternative to SFT for the development of LLMs, which efficiently fine-tune the model with merely dozens to thousands of samples to excel at domain-specific tasks. It uses predefined task-specific reward functions, eliminating the need for costly human annotations. Recent works, such as DeepSeek-R1 (Guo et al., 2025), demonstrate the effectiveness of rule-based RL in mathematical problem solving by evaluating the correctness of the solution, while others (Liu et al., 2025; Wang et al., 2025; Peng et al., 2025; Chen et al., 2025) extend the algorithm to multimodal models, achieving notable improvements in vision-related tasks such as image grounding and object detection. By focusing on measurable objectives, rule-based RL enables practical and versatile model optimization across both textual and multimodal domains, offering significant advantages in terms of efficiency, scalability, and reduced reliance on large datasets.
Previous studies targeting traditional vision-related tasks always rely on the traditional Intersection over Union (IoU) metric commonly used for grounding and detection tasks. In this work, we extend the rule-based RL paradigm to a new application domain by focusing on GUI action prediction tasks driven by low-level instructions. To achieve this, MLLM generates multiple responses (trajectories) that contain the reasoning tokens and the final answers for each input. Then our proposed reward function evaluates each response and updates the model by policy optimization, such as GRPO (Shao et al., 2024), to improve its reasoning ability. Our reward function contains the action type reward, the action argument reward, along with the common format reward. In detail, (1) the action type reward is determined by whether the predicted action type matches the ground truth; (2) the action argument reward (focused on Click), is evaluated by whether the predicted coordinates fall within the ground truth bounding box; (3) the format reward is evaluated by whether the model provides both the reasoning process and the final answer. This flexible and effective reward mechanism is well aligned with the objectives of general GUI-related tasks, ensuring both accuracy and interpretability in the model’s performance.
Regarding data preparation, we follow Muennighoff et al. (2025) and select just 130+ training mobile samples according to three criterion: difficulty, diversity, and quality, making our method remarkably data-efficient. Experiments demonstrate that UI-R1 achieves significant performance improvements on out-of-domain (OOD) data, including entries from desktop and web platforms, indicating the potential of rule-based RL to tackle complex GUI-related tasks across diverse domains effectively.
@@
  • We propose UI-R1, which enhances MLLM’s reasoning capabilities on GUI action prediction tasks through DeepSeek R1 style reinforcement learning.
  • We design a unified rule-based action reward function that effectively aligns with the objectives of common GUI tasks.
  • We utilize the three-stage data selection method and collect only 130+ high-quality training data from the mobile domain. Despite limited data, our model UI-R1-3B achieves notable performance gains on out-of-domain benchmarks, such as those from desktop and web platforms, showcasing adaptability and generalization capability in GUI-related tasks.
@@ 2.1 GUI Agents
Starting with CogAgent (Hong et al., 2024), researchers have used MLLMs for GUI-related tasks, including device control, task completion, GUI understanding, and more. One line of work, such as the AppAgent series (Zhang et al., 2023; Li et al., 2024b) and the Mobile-Agent series (Wang et al., 2024b; a), integrates commercial generalist models like GPT for planning and prediction tasks. These agents rely heavily on prompt engineering and multi-agent collaboration to execute complex tasks, making them adaptable but dependent on careful manual design for optimal performance. Another branch of research focuses on fine-tuning smaller open-source MLLMs on task-specific GUI datasets (Rawles et al., 2023; Li et al., 2024a; Chai et al., 2024; Gou et al., 2024) to create specialist agents. For example,  Chai et al. (2024) enhances agents by incorporating additional functionalities of the GUI element in the Android system, while UGround(Gou et al., 2024) develops a special GUI grounding model tailored for precise GUI element localization.  Wu et al. (2024) develops a foundational model for GUI action prediction. Moving beyond task-specific fine-tuning, UI-TARs (Qin et al., 2025) introduces a more comprehensive approach by combining GUI-related pretraining with task-wise reasoning fine-tuning, aiming to better align models with the intricacies of GUI interactions. Despite their differences, all of these existing agents share a common reliance on the SFT paradigm. This training approach, while effective, depends heavily on large-scale, high-quality labeled datasets.
@@ 2.2 Rule-Based Reinforcement Learning
Rule-based reinforcement learning (RL) has recently emerged as an efficient alternative to traditional training paradigms by leveraging predefined rule-based reward functions to guide model behavior. DeepSeek-R1 (Guo et al., 2025) first introduced this approach, using reward functions based on predefined criteria, such as checking whether an LLM’s final answer matches the ground truth for math problems. The reward focuses solely on the final results, leaving the reasoning process to be learned by the model itself.  Zeng et al. (2025) reproduces the algorithm on models with smaller sizes and illustrates its effectiveness on small language models. Subsequent works (Chen et al., 2025; Shen et al., 2025; Liu et al., 2025; Wang et al., 2025; Peng et al., 2025; Meng et al., 2025), extended the paradigm to multimodal models by designing task-specific rewards for visual tasks, including correct class predictions for image classification and IoU metrics for image grounding and detection. These studies demonstrate the adaptability of rule-based RL for both pure-language and multimodal models. By focusing on task-specific objectives without requiring extensive labeled datasets or human feedback, rule-based RL shows strong potential as a scalable and effective training paradigm across diverse tasks.
@@ 3 Method
UI-R1 is a reinforcement learning (RL) training paradigm designed to enhance a GUI agent’s ability to successfully complete low-level instructional tasks. We define “low-level instructions” as directives that guide the agent to perform actions based on a single state (e.g., a GUI screenshot), consistent with the definition in AndroidControl (Li et al., 2024a). For example, “Click the menu icon in the top left corner” represents a low-level instruction, whereas “Create an event for 2 PM tomorrow” is a high-level instruction. The specifics of the training data selection and reward function design are detailed in the following sections. Figure 2 illustrates the main parts of the framework.
@@ 3.1 Preliminary
Many rule-based RL works (Guo et al., 2025; Zeng et al., 2025; Liu et al., 2025) adopt the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024) for RL training. GRPO offers an alternative to commonly used Proximal Policy Optimization (PPO) (Schulman et al., 2017) by eliminating the need for a critic model. Instead, GRPO directly compares a group of candidate responses to determine their relative quality.
@@ 3.2 Rule-Based Action Rewards
The rule-based reward function introduced by DeepSeek-R1 (Guo et al., 2025) represents a foundational step in rule-based RL by simply evaluating whether model predictions exactly match ground-truth answers. This straightforward approach efficiently aligns models with preference alignment algorithms and provides clear optimization signals. For vision-related tasks, works such as VLM-R1 (Shen et al., 2025) and Visual-RFT (Liu et al., 2025) extend this idea by designing task-specific rewards. For image grounding tasks, they compute the IoU between the predicted and ground-truth bounding boxes as the reward. Similarly, for image classification tasks, rewards are determined by checking whether the predicted and ground-truth classes match.
In GUI-related tasks, the ability of GUI grounding and understanding is a critical requirement for agents. Unlike traditional image grounding tasks, GUI grounding requires agents to identify where specific actions, such as click, should be performed on a given GUI screenshot. To address this unique gap, we propose a reward function tailored for GUI tasks, as defined in Equation 2:
$R_{\mathcal{A}}=R_{\mathcal{T}}+R_{\mathcal{C}}+R_{\mathcal{F}},$ (2)
where the predicted action $\mathcal{A}=\{\mathcal{T},\mathcal{C}\}$ consists of two components: $\mathcal{T}$, which represents the action type (e.g., click, swipe), and $\mathcal{C}$, which represents the click coordinate. $R_{\mathcal{F}}$ represents the common response format reward.
@@
During training, we incorporate the widely-used format reward to guide the model in generating its reasoning process and final answer in a structured format. This decision is based on our simple experiment that agents producing reasoning processes outperform those directly outputting action predictions by approximately 6%. The reasoning process plays a key role in the model’s self-learning and iterative improvement during reinforcement fine-tuning, while the reward tied to the final answer drives optimization. The format reward, denoted as $\boldsymbol{R_{\mathcal{F}}}$, ensures that the model’s predictions follow the required HTML tag format, specifically using $<$$\boldsymbol{think}$$>$ for the reasoning process and $<$$\boldsymbol{answer}$$>$ for the final answer. This structured output not only enhances clarity, but also ensures consistency in the model’s predictions.
@@ 3.3 Training Data Selection
Compared to SFT, rule-based RL has demonstrated the capability to achieve comparable or even superior performance on mathematical and vision-related tasks using only a limited number of training samples (Zeng et al., 2025; Liu et al., 2025). Building on this efficiency and inspired by s1 (Muennighoff et al., 2025), we implement a three-stage data selection process to refine open-source GUI-related datasets based on three key principles: Quality, Difficulty, and Diversity. The detailed distribution of the dataset can be found in Appendix A.2.
@@ 4.2 Action Prediction Capability
4.3 Performance Factor Study
Data Scale
@@ 5 Conclusion
We propose the UI-R1 framework, which extends rule-based reinforcement learning to GUI action prediction tasks, offering a scalable alternative to traditional Supervised Fine-Tuning (SFT). We designed a novel reward function that evaluates both action type and arguments, enabling efficient learning with reduced task complexity. Using only 130+ training samples from the mobile domain, our approach achieves significant performance improvements and strong generalization to out-of-domain datasets, including desktop and web platforms. The results demonstrate the adaptability, data efficiency, and ability of the rule-based RL to handle specialized tasks effectively.

### v2

@@ Abstract
The recent DeepSeek-R1 has showcased the emergence of reasoning capabilities in LLMs through reinforcement learning (RL) with rule-based rewards. Despite its success in language models, its application in multimodal domains, particularly in graphic user interface (GUI) agent tasks, remains under-explored. To address this issue, we propose UI-R1, the first framework to explore how rule-based RL can enhance the reasoning capabilities of multimodal large language models (MLLMs) for GUI action prediction tasks. Specifically, UI-R1 introduces a novel rule-based action reward, enabling model optimization via policy-based algorithms such as Group Relative Policy Optimization (GRPO). For efficient training, we curate a small yet high-quality dataset of 136 challenging tasks, encompassing five common action types on mobile devices. Experimental results demonstrate that our proposed UI-R1-3B achieves significant improvements over the base model (i.e. Qwen2.5-VL-3B) on both in-domain (ID) and out-of-domain (OOD) tasks, with average accuracy gains of 22.1% on ScreenSpot, 6.0% on ScreenSpot-Pro, and 12.7% on AndroidControl. Furthermore, UI-R1-3B delivers competitive performance compared to larger models (e.g., OS-Atlas-7B) trained via supervised fine-tuning (SFT) on 76K samples. These results underscore the potential of rule-based reinforcement learning to advance GUI understanding and control, paving the way for future research in this domain. Code websit: https://github.com/lll6gg/UI-R1.
@@ 1 Introduction
Rule-based reinforcement learning or reinforcement fine-tuning (RFT) has recently emerged as an efficient and scalable alternative to SFT for the development of LLMs, which efficiently fine-tune the model with merely dozens to thousands of samples to excel at domain-specific tasks. It uses predefined task-specific reward functions, eliminating the need for costly human annotations. Recent works, such as DeepSeek-R1 (Guo et al., 2025), demonstrate the effectiveness of rule-based RL in mathematical problem solving by evaluating the correctness of the solution, while others (Liu et al., 2025b; Wang et al., 2025; Peng et al., 2025; Chen et al., 2025a; Huang et al., 2025; Zhou et al., 2025; Chen et al., 2025b) extend the algorithm to multimodal models, achieving notable improvements in vision-related tasks such as image grounding and object detection. By focusing on measurable objectives, rule-based RL enables practical and versatile model optimization across both textual and multimodal domains, offering significant advantages in terms of efficiency, scalability, and reduced reliance on large datasets.
However, existing related studies always target on general vision-related tasks like grounding and detection using Intersection over Union (IoU) metric. In this work, we extend the rule-based RL paradigm to a new application domain by focusing on GUI action prediction tasks driven by low-level instructions. To achieve this, MLLM generates multiple responses (trajectories) that contain the reasoning tokens and the final answers for each input. Then our proposed reward function evaluates each response and updates the model by policy optimization, such as GRPO (Shao et al., 2024). In detail, our action-based reward function contains the action type reward, the action argument reward, along with the commonly used format reward. This flexible and effective reward mechanism is well aligned with the objectives of general GUI-related tasks, enhancing model’s reasoning capabilities of action prediction by iteratively self-learning.
Regarding data preparation, we follow Muennighoff et al. (2025) and select just 130+ training mobile samples according to three criterion: difficulty, diversity, and quality, making our method remarkably data-efficient. Experiments demonstrate that UI-R1 achieves significant performance improvements on out-of-domain gounding tasks like ScreenSpot-Pro (Li et al., 2025) and computer scenarios in ScreenSpot (Cheng et al., 2024), indicating the potential of rule-based RL to tackle complex GUI-related tasks across diverse domains effectively.
@@
  • We propose UI-R1, the first framework which enhances MLLM’s reasoning capabilities on GUI action prediction tasks through DeepSeek R1 style reinforcement learning. We believe our exploration can inspire further advancements in the field.
  • We design a rule-based action reward function that effectively aligns with the objectives of common GUI tasks, facilitating the self-refinement and iterative optimization of the policy model. We also performed some ablation studies to demonstrate its efficiency and rationality.
  • We utilize the three-stage data selection method and collect only 130+ high-quality training data from the mobile domain. Despite limited data, our proposed UI-R1-3B achieves notable performance gains on out-of-domain benchmarks, such as those from desktop and web platforms, showcasing adaptability and generalization capability in GUI-related tasks.
@@ 2.1 GUI Agents
Starting with CogAgent (Hong et al., 2024), researchers have used MLLMs for GUI-related tasks, including device control, task completion, GUI understanding, and more (Liu et al., 2025a). One line of work, such as the AppAgent series (Zhang et al., 2023; Li et al., 2024b) and the Mobile-Agent series (Wang et al., 2024b; a), integrates commercial generalist models like GPT for planning and prediction tasks. These agents rely heavily on prompt engineering and multi-agent collaboration to execute complex tasks, making them adaptable but dependent on careful manual design for optimal performance. Another branch of research focuses on fine-tuning smaller open-source MLLMs on task-specific GUI datasets (Rawles et al., 2023; Li et al., 2024a; Chai et al., 2024; Gou et al., 2024) to create specialist agents. For example,  Chai et al. (2024) enhances agents by incorporating additional functionalities of the GUI element in the Android system, while UGround(Gou et al., 2024) develops a special GUI grounding model tailored for precise localization of the GUI element.  Wu et al. (2024) develops a foundational model for GUI action prediction. Moving beyond task-specific fine-tuning, UI-TARs (Qin et al., 2025) introduces a more comprehensive approach by combining GUI-related pretraining with task-wise reasoning fine-tuning, aiming to better align models with the intricacies of GUI interactions. Despite their differences, all of these existing agents share a common reliance on the SFT paradigm. This training approach, while effective, depends heavily on large-scale, high-quality labeled datasets.
@@ 2.2 Rule-Based Reinforcement Learning
Rule-based reinforcement learning has recently emerged as an efficient alternative to traditional training paradigms by leveraging predefined rule-based reward functions to guide model behavior. DeepSeek-R1 (Guo et al., 2025) first introduced this approach, using reward functions based on predefined criteria, such as checking whether an LLM’s final answer matches the ground truth for math problems. The reward focuses solely on the final results, leaving the reasoning process to be learned by the model itself.  Zeng et al. (2025) reproduces the algorithm on models with smaller sizes and illustrates its effectiveness on small language models. Subsequent works (Chen et al., 2025a; Shen et al., 2025; Liu et al., 2025b; Wang et al., 2025; Peng et al., 2025; Meng et al., 2025), extended the paradigm to multimodal models by designing task-specific rewards for visual tasks, including correct class predictions for image classification and IoU metrics for image grounding and detection. These studies demonstrate the adaptability of rule-based RL for both pure-language and multimodal models. By focusing on task-specific objectives without requiring extensive labeled datasets or human feedback, rule-based RL shows strong potential as a scalable and effective training paradigm across diverse tasks.
@@ 3 Method
UI-R1 is a reinforcement learning training paradigm designed to enhance a GUI agent’s ability to successfully complete low-level instructional tasks. We define “low-level instructions” as directives that guide the agent to perform actions based on a single state (e.g., a GUI screenshot), consistent with the definition in AndroidControl (Li et al., 2024a). For example, “Click the menu icon in the top left corner” represents a low-level instruction, whereas “Create an event for 2 PM tomorrow” is a high-level instruction. The specifics of the training data selection and reward function design are detailed in the following sections. Figure 2 illustrates the main parts of the framework.
@@ 3.1 Preliminary
Many rule-based RL works (Guo et al., 2025; Zeng et al., 2025; Liu et al., 2025b) adopt the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024) for RL training. GRPO offers an alternative to commonly used Proximal Policy Optimization (PPO) (Schulman et al., 2017) by eliminating the need for a critic model. Instead, GRPO directly compares a group of candidate responses to determine their relative quality.
@@ 3.2 Rule-Based Action Rewards
The rule-based reward function introduced by DeepSeek-R1 (Guo et al., 2025) represents a foundational step in rule-based RL by simply evaluating whether model predictions exactly match ground-truth answers. This straightforward approach efficiently aligns models with preference alignment algorithms and provides clear optimization signals. For vision-related tasks, works such as VLM-R1 (Shen et al., 2025) and Visual-RFT (Liu et al., 2025b) extend this idea by designing task-specific rewards. For image grounding tasks, they compute the IoU between the predicted and ground-truth bounding boxes as the reward. Similarly, for image classification tasks, rewards are determined by checking whether the predicted and ground-truth classes match.
In GUI-related tasks, the ability to ground and understand the GUI is a critical requirement for agents. Unlike traditional image grounding tasks, GUI grounding requires agents to identify where specific actions, such as click, should be performed on a given GUI screenshot. To address this unique gap, we propose a reward function tailored for GUI tasks, as defined in Equation 2:
$R=R_{\mathcal{T}}+R_{\mathcal{C}}+R_{\mathcal{F}},$ (2)
where the predicted action $\mathcal{A}=\{\mathcal{T},\mathcal{C}\}$ consists of two components: $\mathcal{T}$, which represents the action type (e.g., click, scroll), and $\mathcal{C}$, which represents the click coordinate. $R_{\mathcal{F}}$ represents the commonly used response format reward.
@@
During training, we incorporate the widely-used format reward to guide the model in generating its reasoning process and final answer in a structured format. This decision is based on our simple experiment that agents producing reasoning processes outperform those directly outputting action predictions by approximately 6% (shown in Appendix B.2). The reasoning process plays a key role in the model’s self-learning and iterative improvement during reinforcement fine-tuning, while the reward tied to the final answer drives optimization. The format reward, denoted as $\boldsymbol{R_{\mathcal{F}}}$, ensures that the model’s predictions follow the required HTML tag format, specifically using $<$$\boldsymbol{think}$$>$ for the reasoning process and $<$$\boldsymbol{answer}$$>$ for the final answer. This structured output not only enhances clarity, but also ensures consistency in the model’s predictions.
@@ 3.3 Training Data Selection
Compared to SFT, rule-based RL has demonstrated the capability to achieve comparable or even superior performance on mathematical and vision-related tasks using only a limited number of training samples (Zeng et al., 2025; Liu et al., 2025b). Building on this efficiency and inspired by s1 (Muennighoff et al., 2025), we implement a three-stage data selection process to refine open-source GUI-related datasets based on three key principles: Quality, Difficulty, and Diversity. The detailed distribution of the dataset can be found in Appendix A.2.
@@ 4.2 Action Prediction Capability
4.3 Key Factor Study
Data Size
@@ 5 Conclusion
We propose the UI-R1 framework, which extends rule-based reinforcement learning to GUI action prediction tasks, offering a scalable alternative to traditional Supervised Fine-Tuning (SFT). We designed a novel reward function that evaluates both action type and arguments, enabling efficient learning with reduced task complexity. Using only 130+ training samples from the mobile domain, our proposed UI-R1-3B achieves significant performance improvements and strong generalization to out-of-domain datasets, including desktop and web platforms. The results demonstrate the adaptability, data efficiency, and ability of the rule-based RL to handle specialized tasks effectively.
@@ B.1 ScreenSpot-V2
B.2 Reasoning
The reasoning capability of the base model (i.e., Qwen2.5-VL-3B) is central to our approach, as it leverages the model’s self-learning and iterative improvement through reinforcement fine-tuning. To enhance the model’s reasoning ability and prediction accuracy, we incorporate reasoning tags into the prompt. To evaluate this, we assess the reasoning performance of Qwen2.5-VL-3B on the ScreenSpot task in Figure 6.
Figure 6: Qwen2.5-VL-3B’s reasoning capability on ScreenSpot.